OpenStudy (ovidia):

-12x-{-16x+[-2x-(14x-11)]} write the terms from the highest to the lowest power of the variable. @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

-12x-{-16x+[-2x-(14x-11)]}=-12x+16x-[-2x-14x+11]=4x-[-16x+11]=4x+16x-11=20x-11

OpenStudy (ovidia):

20x-11? @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

right

OpenStudy (ovidia):

can you help me with another one?

OpenStudy (ovidia):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

ok... but you are not trying.. you try. i will help then.. simple formula -(a+b)=-a-b -(a-b)=-a+b.